### v3

# GreenHyperSpectra: A multi-source hyperspectral dataset for global vegetation trait prediction

Abstract
Plant traits such as leaf carbon content and leaf mass are essential variables in the study of biodiversity and climate change. However, conventional field sampling cannot feasibly cover trait variation at ecologically meaningful spatial scales. Machine learning represents a valuable solution for plant trait prediction across ecosystems, leveraging hyperspectral data from remote sensing. Nevertheless, trait prediction from hyperspectral data is challenged by label scarcity and substantial domain shifts (e.g. across sensors, ecological distributions), requiring robust cross-domain methods. Here, we present GreenHyperSpectra, a pretraining dataset encompassing real-world cross-sensor and cross-ecosystem samples designed to benchmark trait prediction with semi- and self-supervised methods. We adopt an evaluation framework encompassing in-distribution and out-of-distribution scenarios. We successfully leverage GreenHyperSpectra to pretrain label-efficient multi-output regression models that outperform the state-of-the-art supervised baseline. Our empirical analyses demonstrate substantial improvements in learning spectral representations for trait prediction, establishing a comprehensive methodological framework to catalyze research at the intersection of representation learning and plant functional traits assessment. We also share the dataset [^1] GreenHyperSpectra dataset: https://huggingface.co/datasets/Avatarr05/GreenHyperSpectra, code and pretrained model objects for this study here.
1 Introduction
Plant functional traits are a fundamental component of biodiversity assessment, offering insights into plant productivity, ecological interactions, resilience, and adaptation to environmental change (cavender2020remote; funk2017revisiting; skidmore2021priority; yan2023essential). Leaf traits such as leaf mass per area, as well as chlorophyll, nitrogen, and carbon content, are key to understanding plant growth dynamics and ecosystem processes such as carbon cycling and productivity (berger2022multi; bongers2021functional; damm2018remote; roscher2012using; zarco2018previsual; zarco2019chlorophyll). The monitoring of these traits is thus crucial for understanding ecosystem function and guiding biodiversity conservation strategies (cavender2022integrating; pettorelli2021time; xu2021ensuring).
Figure 1: Overview of the semi/self-supervised framework for multi-trait regression task.
Initiatives such as the Intergovernmental Science-Policy Platform on Biodiversity and Ecosystem Services (IPBES; (diaz2015ipbes; ipbes2019global)) have raised global awareness around the urgent need for monitoring functional traits and their diversity across spatial scales. However, we still lack efficient tools to track these functional traits in space and time. Hyperspectral remote sensing from airborne and satellite systems has emerged as a promising tool to bridge this gap, enabling non-destructive, scalable, and repeatable reflectance measurements that can be used to predict these traits (berger2020crop; cherif2023spectra; jetz2016monitoring; serbin2019arctic; van2018functional). Hyperspectral sensors measure radiation reflected from the ground across hundreds of narrow, contiguous spectral bands, spanning the visible to shortwave infrared (VNIR + SWIR) domains. These measurements are informative for trait prediction, as they are directly influenced by the chemical and structural characteristics of plant leaves and canopies (jacquemoud2019leaf). Plant trait prediction from hyperspectral data is inherently a regression problem, and was initially often explored with Partial Least Squares Regression (PLSR, (geladi1986partial)) to link hyperspectral observations to individual traits (feilhauer2010brightness; serbin2015remotely; helsen2021evaluating; ustin2009retrieval). However, non-parametric machine learning methods, in particular deep learning, have recently been explored to offer greater flexibility in modeling complex, non-linear trait-spectral relationships and trait-trait interactions (cherif2023spectra; tsakiridis2020simultaneous; he2016novel). In this context, trait prediction is framed as a multi-output regression task within a multi-task learning framework, where the outputs, representing multiple plant traits, are inherently correlated.
Trait prediction poses fundamental machine learning challenges including heterogeneous target distributions requiring specialized multi-task methods cherif2023spectra, extreme label scarcity, and significant distributional biases (e.g. spatial, Figure 2). Current supervised approaches demonstrate limited cross-domain generalization due to training on sparse, non-representative datasets (danilevicz2022plant; pichler2020machine). Hyperspectral data further compound these challenges exhibiting substantial covariate shifts across acquisition conditions, sensor configurations and resolutions, radiometric calibrations and variable input modalities. To address these limitations, we introduce GreenHyperSpectra, a large spectral dataset designed to improve representation robustness against domain adaptation challenges while being collected from multiple ecosystems, instruments, spatial resolutions, and acquisition conditions. This dataset enables semi- and self-supervised learning applications, which take advantage of vast unlabeled spectral data, providing extensive coverage and variability to facilitate benchmarking.
Our contributions include: 1 Building GreenHyperSpectra, a dataset for pretraining consisting of cross-domain samples and substantially expanding available datasets for representation learning; 2 framing a suite of semi- and self-supervised methods for multi-output regression with one dimensional (1D) hyperspectral data; 3 comparing these methods with fully supervised baseline, highlighting the superior performance of the former, particularly in scenarios with limited labeled data; 4 Testing how well such methods generalize across variable inputs, representing the diversity of sensor settings (full-range (VNIR+SWIR) vs. half-range (VNIR-only), see Figure 1).
2 Related work
Despite increasing hyperspectral data availability, plant trait prediction is still largely constrained by the lack of large annotated datasets (danilevicz2022plant; todman2023small). Trait labels are costly and time-consuming to obtain, often requiring field sampling and laboratory analysis  (cornelissen2003handbook; baraloto2010functional). Available datasets lack harmonization and often differ in sampling strategy, measurement assumptions and protocols. As a result, most labeled datasets are geographically and ecologically limited, with sparse coverage across space and time, as well as across ecosystems and acquisition conditions. To address this, previous studies explored synthetic datasets generated from Radiative Transfer Models (RTMs, (feret2008prospect)) that simulate canopy spectral responses under diverse conditions. In this context, it motivated hybrid approaches that combine RTM-based simulations with machine learning (danner2021efficient; mederer2025plant; parigi2024towards; tagliabue2022hybrid; verrelst2015optical; wang2021airborne).
Figure 2: Spatial coverage of the datasets. Points represent sample locations of GreenHyperSpectra compared to the existing labeled dataset. GreenHyperSpectra data span diverse vegetation type and acquisition conditions.
However, several comparisons reveal important limitations, with models trained on real, multi-site datasets consistently outperforming those using only synthetic spectra (mederer2025plant; parigi2024towards), highlighting a persistent domain gap between simulated and field data.
In this regard, there is a need of large-scale unlabeled spectral datasets from real measurements to pretrain models. To support this, a number of benchmarking efforts have been introduced (fuchs2023hyspecnet; wang2025hypersigma; braham2024spectralearth). However, most of these initiatives are built from a single sensor type, restricting spectral diversity and limiting generalization to new acquisition conditions. Moreover, these datasets often consist of full hyperspectral imagery, where spatially contiguous pixels are subject to spatial autocorrelation. This spatial redundancy reduces the spectral variability necessary for training models that generalize well in trait prediction task, which depend on spectral rather than spatial information (thoreau2024toulouse). They also typically include a broad mix of land cover types, including non-vegetated surfaces. This introduces inefficiencies, as vegetated pixels must be sampled through additional preprocessing.
Large-scale unlabeled datasets have been leveraged for self-supervised learning to improve natural image representations Chen_2021_ICCV; Assran2023Self; oquab2024dinov as well as in remote sensing with multispectral data tseng2023lightweight; NEURIPS2023_11822e84; xiong2024neuralplasticityinspiredmultimodalfoundation; szwarcman2025prithvieo20versatilemultitemporalfoundation; bountos_fomo_2025; astruc2025anysatearthobservationmodel; tseng2025galileolearninggloballocal. Semi- and self-supervised learning techniques are increasingly being explored to exploit hyperspectral data for image classification, segmentation, and super-resolution (ahmad2021hyperspectral; ranjan2023unlocking; wang2022self). These methods are designed to better learn the spectral representation of hyperspectral remote sensing data to reduce the need for reference labels. Approaches such as masked autoencoders (MAE) (cong2022satmae; hong2021spectralformer; reed2023scale; scheibenreif2023masked; thoreau2024toulouse), contrastive learning frameworks (chen2020simple; fuller2023croma; he2020momentum), generative networks (GAN) (alipour2020structure; he2017generative; kwak2023semi; roy2021generative; zhan2018semi; zhu2018generative), and autoencoders (AE) (ahmad2019segmented; gallo2023self; thoreau2024toulouse; zhao2017spectral) have been successfully applied for land cover classification. Whereas most attempts at trait prediction using hyperspectral data relied on fully supervised pipelines, notably PLSR (serbin2015remotely; singh2015imaging; ustin2009retrieval; wang2020foliar), Gaussian Process Regression (GPR) (tagliabue2022hybrid; verrelst2013gaussian; verrelst2012gaussian), ANN (schlerf2006inversion) and deep learning methods (cherif2023spectra; pullanagari2021field; shi2022convolution), they are cardinally constrained by label scarcity, hampering their ability to reliably generalize across ecosystems, sensor platforms, and acquisition conditions. Applying semi- and self-supervised methods to trait prediction remains largely unexplored and offers a compelling direction for investigation. Recent semi- and self-supervised applications in trait prediction include vision transformers for nitrogen estimation from simulated data (gallo2023self) and Long short-term memory (LSTM) models for chlorophyll prediction with limited spectral bands (zhao2023improving) (i.e. VNIR-only). While semi- and self-supervised methods show promise for trait prediction, existing models remain constrained to single traits. Moreover, existing models are typically limited to specific sensor configurations and experimental conditions, limiting generalization across sensor modalities (e.g. full-range vs. half-range spectrometers), acquisition geometries, and vegetation types. This underscores the need for flexible approaches that can handle heterogeneous inputs while supporting transferable predictions in diverse real-world ecological scenarios. To the best of our knowledge, no semi- or self-supervised method addresses trait prediction via multi-output regression.
3 The GreenHyperSpectra dataset
We introduce a large-scale, multi-source hyperspectral dataset comprising over $140,000$ vegetation canopy surface reflectance spectra captured across diverse continents, ecosystems, sensor platforms, spatial resolutions, and measurement geometries. Unlike existing benchmarks of hyperspectral data limited to single sensors or narrow ecological domains, our dataset features a substantially larger pretraining spectral dataset supporting semi- and self-supervised learning approaches.
Acquisition platforms and sensor diversity.
We curated spectral data from multiple instruments across three primary platforms: proximal, airborne, and spaceborne (Figure 2). All data were processed to the level of at-surface reflectance. Proximal measurements were obtained using field spectrometers such as the ASD FieldSpec and SVC HR-1024i, typically positioned in a close range in nadir orientation to record top-of-canopy reflectance. Airborne data were acquired using high-spectral resolution sensors, including the AVIRIS-Next Generation, AVIRIS-Classic, NEON Airborne Observation Platform (AOP) and Specim AISAFenix instruments, which cover landscape-level vegetation scenes with variable viewing geometries and meter-scale spatial resolutions. Spaceborne acquisitions were collected from missions such as PRISMA, Hyperion, EMIT, and EnMAP, offering a larger scale
Platform GSD Spectral res. #Samples
Proximal <1 m 1–4 nm 5620
Airborne 1–20 m 3–7 nm 96699
Spaceborne 30–60 m 6–12 nm 36059
Table 1: Specifications of spectroscopy instruments with different platforms.
observations at 30-60 m resolution with varying viewing geometry. Table 1 summarizes the platforms, spectral properties, and scene-level characteristics associated with each acquisition (more details see Appendix A). The pre-processing of spectra harmoinization is described in Appendix  A.
The multi-platform nature of our dataset introduces valuable reflectance signal variability through differences in spatial and spectral resolution, sun-sensor geometry, scene heterogeneity, background conditions and pre-processing from radiance to reflectance. This variability, often lacking in single-platform or synthetic datasets, is essential to develop generalizable models capable of scaling across diverse remote sensing contexts (cherif2023spectra). While satellite-based datasets such as SpectralEarth (braham2024spectralearth) provide temporally rich but sensor-specific imagery, our dataset uniquely incorporates multi-sensor observations across spaceborne, airborne, and proximal platforms.
Spatial and temporal coverage.
The dataset includes samples from diverse biomes, with acquisitions spanning from 1992 to 2024, capturing broad ecological and climatic variability across a wide range of environments. Figure 2 maps the global spatial distribution of GreenHyperSpectra and a pool of previously aggregated datasets for plant trait prediction (see details in § 4 ). While the compiled labeled dataset is spatially limited, GreenHyperSpectra encompasses substantially broader spatial coverage and environmental heterogeneity, better representing real-world remote sensing operational conditions (more details in Appendix A).
4 Benchmarking methods and protocols
Trait-annotated dataset.
[…]
Data splitting and evaluation protocol.
To ensure consistent representation of all contributing data sources during training, we divide GreenHyperSpectra into 20 non-overlapping subsets. In each split, the proportion of samples from any given data source matches that dataset’s overall contribution to the full merged dataset. This stratified splitting strategy maintains the natural diversity of vegetation types, sensors, and acquisition conditions, while preventing bias from individual sources by creating consistent and representative subsets suitable for semi- and self-supervised methods. For the labeled dataset, we define standardized train and validation splits using a 80/20 hold-out strategy. The 80% portion is combined with the pretraining spectral dataset for calibration, while the remaining 20% is fixed for all experiments and used to evaluate all methods. For out of distribution (OOD) evaluation experiments (detailed in §5), we perform cross validation across the 50 labeled datasets. Specifically, from the 50 annotated sub-datasets, we hold out five datasets at a time for testing. The remaining datasets are used for training, with their data further split into 80% for training and 20% for validation.
(a) Semi-supervised generative adversarial network framework (SR-GAN).
(b) RTM-based autoencoder framework (RTM-AE).
(c) Masked-autoencoder framework (MAE).
Figure 3: Overview of the semi- and self-supervised frameworks. (3(a)) The semi-supervised regression GAN framework (SR-GAN): the generator maps a random noise $z$ to synthetic samples $\hat{x}$, while the discriminator processes 1 fake samples ($x_{\text{fake}}$), 2 unlabeled real samples ($x_{\text{unlb}}$), and 3 labeled real data samples ($x_{\text{lb}}$) with associated traits (y), optimizing fake ($L_{\text{fake}}$), unlabeled ($L_{\text{unlb}}$), and labeled ($L_{\text{lb}}$) losses respectively. (3(b)) The RTM-based autoencoder (RTM-AE) predicts traits from labeled embeddings while reconstructing spectra ($x\to\hat{x}$, ($L_{\text{recon}}$)). (3(c)). The 1D masked autoencoder framework (1D-MAE) reconstructs masked spectra through tokenization, ($L_{\text{recon}}$); the learned representations are then used for trait prediction ($L_{\text{lb}}$). Abbreviations: $x_{\text{fake}}$: generated fake spectra from the generator; $x_{\text{unlb}}$: unlabeled sample from GreenHyperSpectra; $x_{\text{lb}}$: spectra sample from the labeled data; $L_{\text{unlb}}$: unlabeled loss; $L_{\text{lb}}$: labeled loss; $L_{\text{recon}}$: reconstruction loss; $L_{\text{fake}}$: feature contrasting loss; RTM: radiative transfer model; AE: autoencoder; MAE: masked autoencoder.
Supervised baseline method.
We consider a supervised CNN-based method (cherif2023spectra; mederer2025plant) as a baseline, selected for its state-of-the-art performance in multi-trait plant prediction from a sparse annotated dataset. It is built upon EfficientNet-B0 (tan2019efficientnet) specifically framed for 1D feature extraction. The network employs multi-output regression to simultaneously predict the seven plant traits, a strategy that demonstrably outperforms single-trait modeling approaches (cherif2023spectra; scutari2014multiple).
Semi-supervised regression generative adversarial network (SR-GAN).
We frame the SR-GAN framework (olmschenk2019generalizing) to address hyperspectral plant trait prediction. Our implementation employs a 1D convolutional GAN architecture designed specifically for spectral data processing. In this setup, the generator learns to produce synthetic reflectance spectra, while the discriminator simultaneously performs trait regression and learns discriminative feature representations. The training objective is formulated as a composite loss that encourages the discriminator to pull real spectral samples closer in the feature space, while pushing representations of generated (synthetic) samples further apart. This contrastive learning approach allows the model to leverage unlabeled data by learning informative spectral embeddings. The overall architecture of the SR-GAN framework is illustrated in Figure 3(a) and detailed formulations of the loss components are provided in Appendix B.1.
Radiative transfer model based autoencoder (RTM-AE).
We introduce a version of the autoencoder framework proposed by (she2024spectra), which replaces the decoder with a non-learnable RTM module to reconstruct spectra, thereby integrating physical constraints into the modeling process. Specifically, our implementation employs PROSAIL-PRO (feret2021prospect), constraining the latent space to correspond directly to plant traits. PROSAIL-PRO is an RTM that combines the leaf reflectance model (PROSPECT,(jacquemoud1990prospect)) with the canopy reflectance model (4SAIL,(verhoef2007unified)). PROSPECT simulates leaf reflectance and transmittance based on biochemical composition and internal structure, while 4SAIL models the propagation of light through a vegetation canopy. Together, they simulate canopy spectral reflectance in the 400–2500 nm range using inputs such as chlorophyll content, leaf area index, and leaf angle. As previously mentioned regarding the gap between RTM-simulated and real-world spectra (§ 2), we address the inherent discrepancies between RTM-generated and observed spectra, primarily resulting from simplified geometric assumptions within the model, by implementing a learnable correction layer that refines the simulated output (she2024spectra). Our enhanced framework introduces three key improvements over the original design (she2024spectra): (1) incorporation of PROSAIL-PRO, (2) application of a supervised loss component targeting trait predictions, and (3) implementation of a composite reconstruction loss combining cosine similarity and mean absolute error to capture both spectral shape characteristics and amplitude information. The overall architecture is illustrated in Figure 3(b) and the specifications are detailed in Appendix B.2 .
Masked autoencoder (MAE).
We adopt a MAE framework, originally designed for land cover classification (thoreau2024toulouse), to predict plant traits with hyperspectral data. The model leverages self-supervised learning by reconstructing randomly masked spectral regions, enabling the extraction of meaningful representations from unlabeled hyperspectral signatures. Similarly to the RTM-AE, our adaptation incorporates a modified reconstruction objective that combines cosine similarity and mean squared error (MSE) with appropriate weighting, allowing the model to capture both spectral shape characteristics and amplitude information. For downstream trait prediction, we attach a multi-output regression head to the latent features and fine-tune the model using labeled data. The overall architecture is illustrated in Figure 3(c) and ablation studies for the MAE architecture are provided in Appendix B.3.
Figure 4: Evaluation of trait prediction with variable-size labeled sets. Validation performance ($R^{2}$) as a function of labeled data percentage used for training. The average $R^{2}$ performance across all traits is indicated by the dashed box. The higher $R^{2}$, the better. For trait abbreviations, see Sec. 4.
5 Experimental settings
This section describes experimental setups used to benchmark and evaluate the performance of models trained on semi- and self-supervised learning fashion for multi-trait plant prediction detailed in § 4. Our experimental framework is structured into four principal components to test the capabilities of models across a range of scenarios reflecting critical use cases: comprehensive benchmarking using full-range (FR) spectra, benchmarking with half-range (HR) spectra, and assessment of OOD generalization capabilities, along with an ablation study on the design of the MAE models. Throughout these experiments, we maintain standardized data splits for both labeled and unlabeled datasets as described in § 4. Complete specifications regarding model hyperparameters, optimization settings, and implementation details are provided in Appendix B.
Full-range trait prediction.
We assess all benchmark models using the full-range spectra spanning 400–2450 nm (1721 bands), encompassing visible through shortwave infrared wavelengths. Results are presented in § 6 with Table 2.
Sample sensitivity analysis.
We examine the impact of label availability by simulating different levels of supervision and varying the amount of labeled data used for training from 20% to 100% while maintaining a consistent unlabeled dataset. Complementing this approach, we conduct experiments varying the quantity of unlabeled training data while maintaining fixed labeled data proportions to determine how unlabeled data volume influences model performance; note that in these experiments, we use only a subset of the full GreenHyperSpectra dataset ($80,000$ samples). Results are presented in § 6, with Figure 4 and  5.
Half-range trait prediction.
A common constraint faced with satellite-based Earth observations is that many sensors do not cover the full spectrum. To evaluate model performance in this scenario, we replicate our benchmark procedure using only the half-range spectral subset spanning 400–900nm (500 bands). All models are trained on this spectral subset. Additionally, we implement an evaluation for the MAE architecture, where a model pretrained on full-range spectra is applied to half-range spectra inputs (this is possible only for the MAE models as the masking procedure means that they can accommodate variable input sizes). Results are presented in § 6 and Table 3.
OOD evaluation.
To assess each model’s robustness to real-world distribution shifts, we perform a cross-dataset evaluation as described in § 4. We compute a macro-level performance metric by aggregating predictions across all held-out datasets. This setup reflects practical challenges in ecological monitoring applications, where spectral variability arises from differences in acquisition conditions, sensor platforms, or environmental contexts. Additionally, this approach ensures a broader coverage of trait value ranges, which often remain underrepresented when test sets are randomly sampled. Due to computational constraints, we conduct this evaluation using a single training run. To reduce the sensitivity of $R^{2}$ to unbalanced number of samples across the 50 aggregated datasets, we compute the macro-average over five random subsamples within each dataset, each constrained to the maximum number of 30 samples allowed per set, and report the mean and standard deviation of the resulting metrics. Results are presented in § 6 and Table 4.
Figure 5: Evaluation of trait prediction with variable-size unlabeled sets. Validation performance ($R^{2}$) as a function of the percentage of unlabeled data used for training. The average $R^{2}$ performance is indicated by the dashed box. The higher $R^{2}$, the better. For trait abbreviations, see Sec. 4.
Ablation studies on MAE.
We conduct comprehensive ablation experiments across several dimensions to consistently evaluate design trade-offs of the MAE models in spectral representation learning. First, we explore architectural complexity through a grid search spanning transformer configurations with varying numbers of layers $\{6,8,10\}$ and attention heads $\{4,8,16\}$. We select the configuration demonstrating optimal performance on the downstream trait prediction task for subsequent experiments. Considering this optimal architecture, we investigate alternative loss formulations for spectral reconstruction. Beyond conventional MSE, we examine hybrid approaches which incorporate cosine similarity loss weighted by a coefficient $\alpha\in\{1,0.1,0.01,0\}$ to enhance capture of spectral shape specificity. Finally, we assess the effect of token granularity by varying patch sizes (10, 20, 40, and 430) used during spectral masking and reconstruction in the full-range scenario. These targeted ablations informed the final MAE configuration used in our other benchmark evaluations. All results of the aforementioned ablation studies are presented in Tables 16, 17 and 18 in Appendix B.3.
Evaluation metrics.
For each experimental setting, we report the performance metrics averaged across three random seeds to measure the variability related to stochastic training effects. Our evaluation framework employs two complementary metrics: the coefficient of determination ($R^{2}$) and the normalized root mean square error (nRMSE). The nRMSE (in %) is computed by normalizing the root mean square error by the range of the traits observations (1–99% quantile), providing a scale-invariant measure of prediction error.
6 Results and discussions
Labeled and unlabeled data regimes.
To assess each model’s sensitivity to the quantity of annotated samples, we analyze $R^{2}$ as a function of the proportion of available labeled and unlabeled data, as shown in Figures 4 and 5 respectively. The corresponding trends for nRMSE are presented in Appendix D. We observed that models leveraging unlabeled data through semi- and self-supervised methods consistently outperformed the fully supervised baseline, particularly in low-data regimes (20–40% labeled data). Notably, semi- and self-supervised methods achieved higher average $R^{2}$ and lower nRMSE scores across most traits as labeled data availability decreased (Fig. 4). This demonstrates that access to a large set of unlabeled spectra through GreenHyperSpectra substantially enhances model performance, leading to improved trait prediction accuracy. Interestingly, varying the size of the dataset for pretraining did not substantially impact performance. This suggests that the stratified splitting protocol (§ 4), ensuring consistent coverage across spectral sources, vegetation types, and acquisition conditions, plays a critical role in efficiently exploiting the available unlabeled data, even when subsampled.
cab cw cm LAI cp cbc car anth average
$R^{2}$ $(\uparrow)$
Supervised 0.517 (± 0.012) 0.621 (± 0.009) 0.667 (± 0.005) 0.561 (± 0.011) 0.651 (± 0.002) 0.679 (± 0.006) 0.544 ± (0.018) 0.454 (± 0.020) 0.587 (± 0.010)
SR_GAN 0.574 (± 0.008) 0.572 (± 0.013) 0.669 (± 0.008) 0.538 (± 0.005) 0.558 (± 0.005) 0.704 (± 0.004) 0.578 (± 0.006) 0.541 (± 0.041) 0.592 (± 0.011)
RTM_AE 0.584 (± 0.023) 0.658 (± 0.011) 0.679 (± 0.021) 0.552 (± 0.029) 0.671 (± 0.015) 0.689 (± 0.021) 0.566 (± 0.038) 0.337 (± 0.125) 0.592 (± 0.035)
MAE_FR_LP 0.462 (± 0.002) 0.514 (± 0.009) 0.577 (± 0.010) 0.470 (± 0.007) 0.464 (± 0.007) 0.611 (± 0.008) 0.412 (± 0.002) 0.217 (± 0.015) 0.466 (± 0.008)
MAE_FR_FT 0.515 (± 0.026) 0.634 (± 0.014) 0.716 (± 0.027) 0.615 (± 0.016) 0.676 (± 0.010) 0.727 (± 0.034) 0.649 (± 0.012) 0.598 (± 0.025) 0.641 (± 0.020)
nRMSE $(\downarrow)$
Supervised 17.341 (± 0.221) 13.103 (± 0.150) 10.671 (± 0.083) 17.487 (± 0.218) 10.229 (± 0.021) 10.634 (± 0.099) 13.647 (± 0.274) 16.465 (± 0.333) 13.697 (± 0.175)
SR_GAN 16.277 (± 0.161) 13.918 (± 0.208) 10.658 (± 0.127) 17.829 (± 0.091) 11.565 (± 0.074) 10.258 (± 0.076) 13.123 (± 0.095) 15.084 (± 0.662) 13.589 (± 0.187)
RTM_AE 16.096 (± 0.437) 12.443 (± 0.201) 10.492 (± 0.350) 17.556 (± 0.573) 9.989 (± 0.224) 10.495 (± 0.350) 13.297 (± 0.596) 18.091 (± 1.761) 13.557 (± 0.561)
MAE_FR_LP 18.297 (± 0.034) 14.830 (± 0.135) 12.052 (± 0.146) 19.094 (± 0.133) 12.746 (± 0.080) 11.747 (± 0.120) 15.501 (± 0.030) 19.727 (± 0.185) 15.499 (± 0.108)
MAE_FR_FT 17.374 (± 0.456) 12.861 (± 0.256) 9.856 (± 0.462) 16.285 (± 0.333) 9.916 (± 0.155) 9.833 (± 0.599) 11.975 (± 0.202) 14.117 (± 0.438) 12.777 (± 0.363)
Table 2: Evaluation of trait prediction with full-range (FR) samples. Trait-wise performance (mean ± standard deviation) includes $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$ metrics. Competing methods are: fully supervised baseline (‘Supervised’); MAE with full-range training and linear probing (‘MAE_FR_LP’); MAE with full-range training and fine-tuning (‘MAE_FR_FT’); RTM-based autoencoder (‘RTM_AE’); and semi-supervised regression GAN (‘SR-GAN’). In RTM-AE, cbc is not directly predicted but is derived from cm and cp estimates (cm – cp). We bold and underline best and second best scores respectively. Trait abbreviations are detailed in Sec. 4.
cab cw cm LAI cp cbc car anth average
$R^{2}$ $(\uparrow)$
Sup_HR 0.277 (± 0.105) 0.072 (± 0.032) 0.197 (± 0.082) 0.048 (± 0.110) 0.197 (± 0.080) 0.219 (± 0.074) 0.126 (± 0.135) 0.166 (± 0.052) 0.163 (± 0.084)
SR-GAN_HR 0.496 (± 0.017) 0.336 (± 0.006) 0.356 (± 0.011) 0.428 (± 0.010) 0.371 (± 0.010) 0.381 (± 0.008) 0.455 (± 0.015) 0.598 (± 0.020) 0.427 (± 0.012)
RTM-AE_HR 0.582 (± 0.024) 0.450 (± 0.014) 0.472 (± 0.031) 0.541 (± 0.019) 0.546 (± 0.019) 0.471 (± 0.043) 0.491 (± 0.029) 0.538 (± 0.010) 0.511 (± 0.023)
MAE_FR_HR_LP 0.466 (± 0.008) 0.220 (± 0.009) 0.271 (± 0.006) 0.378 (± 0.004) 0.253 (± 0.008) 0.274 (± 0.003) 0.434 (± 0.004) 0.234 (± 0.013) 0.316 (± 0.007)
MAE_FR_HR_FT 0.578 (± 0.011) 0.553 (± 0.009) 0.655 (± 0.012) 0.540 (± 0.012) 0.612 (± 0.022) 0.642 (± 0.009) 0.512 (± 0.018) 0.433 (± 0.032) 0.566 (± 0.015)
MAE_HR_LP 0.493 (± 0.004) 0.221 (± 0.014) 0.247 (± 0.007) 0.435 (± 0.003) 0.280 (± 0.006) 0.279 (± 0.003) 0.375 (± 0.010) 0.376 (± 0.019) 0.338 (± 0.008)
MAE_HR_FT 0.518 (± 0.038) 0.392 (± 0.045) 0.397 (± 0.028) 0.567 (± 0.016) 0.478 (± 0.057) 0.402 (± 0.048) 0.418 (± 0.034) 0.547 (± 0.056) 0.465 (± 0.040)
nRMSE $(\downarrow)$
Sup_HR 21.177 (± 1.575) 20.501 (± 0.333) 16.560 (± 0.831) 25.575 (± 1.465) 15.479 (± 0.757) 16.601 (± 0.774) 18.856 (± 1.492) 20.346 (± 0.634) 19.387 (± 0.982)
SR-GAN_HR 17.706 (± 0.297) 17.339 (± 0.073) 14.869 (± 0.131) 20.007 (± 0.175) 13.809 (± 0.108) 14.827 (± 0.092) 14.914 (± 0.211) 14.133 (± 0.355) 15.950 (± 0.180)
RTM-AE_HR 16.125 (± 0.460) 15.772 (± 0.199) 13.460 (± 0.395) 17.777 (± 0.370) 11.725 (± 0.242) 13.698 (± 0.558) 14.415 (± 0.413) 15.155 (± 0.163) 14.766 (± 0.350)
MAE_FR_HR_LP 17.742 (± 0.140) 17.242 (± 0.102) 16.930 (± 0.065) 19.352 (± 0.056) 18.377 (± 0.095) 17.136 (± 0.040) 15.197 (± 0.052) 19.545 (± 0.165) 17.690 (± 0.089)
MAE_FR_HR_FT 15.771 (± 0.202) 13.055 (± 0.126) 11.652 (± 0.205) 16.647 (± 0.213) 13.242 (± 0.370) 12.037 (± 0.154) 14.109 (± 0.255) 16.805 (± 0.469) 14.165 (± 0.249)
MAE_HR_LP 17.768 (± 0.070) 18.769 (± 0.163) 16.072 (± 0.071) 19.729 (± 0.050) 14.764 (± 0.065) 15.995 (± 0.036) 15.976 (± 0.125) 17.609 (± 0.262) 17.085 (± 0.106)
MAE_HR_FT 17.313 (± 0.692) 16.578 (± 0.621) 14.381 (± 0.337) 17.263 (± 0.314) 12.561 (± 0.692) 14.563 (± 0.589) 15.415 (± 0.453) 14.973 (± 0.948) 15.381 (± 0.581)
Table 3: Evaluation of trait prediction with half-range (HR) samples. Trait-wise performance (mean ± standard deviation) includes $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$ metrics. Competing methods are: supervised baseline with HR settings (‘Sup_HR’); semi-supervised SR-GAN with HR settings (‘SR-GAN_HR’); RTM-based autoencoder with HR settings (‘RTM-AE_HR’); MAE pretrained on full-range and fine-tuned with linear probing (‘MAE_FR_HR_LP’); MAE pretrained on full-range and fine-tuned (‘MAE_FR_HR_FT’); MAE pretrained on HR and fine-tuned with linear probing (‘MAE_HR_LP’); and MAE pretrained on HR and fine-tuned (‘MAE_HR_FT’). In RTM-AE, cbc is not directly predicted but is derived from cm and cp estimates (cm – cp). We bold and underline best and second best scores respectively. Trait abbreviations are detailed in Sec. 4.
Full and half range spectra analyses.
Trait-specific results, reported with $R^{2}$ and nRMSE scores, are summarized in Tables 2 and 3 for the full- and half-range experiments, respectively. Among all competing methods, the fine-tuned MAE (MAE-FR-FT) outperformed all other methods on most traits when trained and tested on full-range spectra, recording the highest $R^{2}$ values and lowest nRMSE scores. Compared to the fully supervised baseline, MAE-FR-FT led to an average improvement of 9% in $R^{2}$ and 6% in nRMSE. These results underscore the effectiveness of MAEs in learning meaningful spectral representations through masked spectral reconstruction. Pretrained MAE models also exhibited strong cross-spectral generalization, performing competitively on half-range data even when pretrained on full-range spectra and applied to half-range data (MAE-FR-HR-FT). It indicates a good feature transferability and adaptability across heterogeneous sensor configurations, particularly valuable for operational deployment with multi-source data streams.
The RTM-AE model, which introduces physical interpretability into the learned latent space, underperformed compared to MAE but consistently achieved the second best results for both full- and half-range experiments. It demonstrates that aligning latent representations with RTMs to enforce physically-constrained embeddings yields promising performance while simultaneously enhancing model explainability through semantically meaningful feature disentanglement and physics-informed representation learning.
To further assess robustness of the approaches under sensor noise such as illumination differences, or sensor-specific signal-to-noise characteristics, we additionally evaluated models’ performances under additive Gaussian noise at inference time (details in Tables 26-30). Zero-mean noise with standard deviations of 0.01, 0.03, and 0.05 was added across all spectral bands. Results show that MAE-FR-FT and RTM-AE are substantially more robust than the supervised baseline and GAN. For instance, at $\sigma=0.05$, MAE-FR-FT retains an $R^{2}$ of 0.331 compared to the baseline dropping to –0.065, and both MAE-FR-FT and RTM-AE exhibit smaller increases in nRMSE under spectral corruption, highlighting their resilience.
In the half-range setting, semi- and self-supervised methods also clearly outperformed the supervised baseline. Gains ranged between 100–200% in $R^{2}$ and 8–27% in nRMSE, reinforcing the value of leveraging spectral variability from GreenHyperSpectra even under reduced spectral coverage (Figures  12 and  13).
cab cw cm LAI cp cbc car anth average
$R^{2}$ $(\uparrow)$
Supervised 0.362 (± 0.048) 0.193 (± 0.053) 0.446 (± 0.049) 0.074 (± 0.031) 0.183 (± 0.041) 0.449 (± 0.052) 0.181 (± 0.045) 0.055 (± 0.079) 0.243 (± 0.050)
SR_GAN 0.300 (± 0.023) 0.350 (± 0.032) 0.507 (± 0.029) -0.199 (± 0.111) 0.273 (± 0.037) 0.548 (± 0.026) 0.221 (± 0.064) 0.197 (± 0.175) 0.275 (± 0.062)
RTM_AE 0.272 (± 0.033) 0.193 (± 0.096) 0.453 (± 0.067) 0.019 (± 0.054) 0.192 (± 0.056) -0.075 (± 0.008) 0.266 (± 0.056) 0.067 (± 0.252) 0.173 (± 0.078)
MAE_FR_LP 0.116 (± 0.028) 0.298 (± 0.028) 0.442 (± 0.039) 0.182 (± 0.059) 0.211 (± 0.032) 0.478 (± 0.044) 0.232 (± 0.020) 0.142 (± 0.153) 0.263 (± 0.050)
MAE_FR_FT 0.271 (± 0.030) 0.28 (± 0.102) 0.575 (± 0.041) 0.229 (± 0.041) 0.275 (± 0.068) 0.582 (± 0.044) 0.165 (± 0.044) 0.112 (± 0.234) 0.311 (± 0.076)
nRMSE $(\downarrow)$
Supervised 19.173 (± 0.695) 25.223 (± 13.109) 14.238 (± 0.496) 22.984 (± 0.475) 17.072 (± 0.669) 14.818 (± 0.560) 19.185 (± 0.581) 23.159 (± 1.644) 19.482 (± 2.278)
SR_GAN 20.098 (± 0.373) 22.394 (± 10.712) 13.445 (± 0.404) 26.075 (± 1.018) 16.108 (± 0.662) 13.438 (± 0.578) 18.698 (± 0.568) 21.360 (± 3.596) 18.952 (± 2.239)
RTM-AE 20.493 (± 0.443) 25.137 (± 13.001) 14.138 (± 0.616) 23.652 (± 0.668) 16.978 (± 0.802) 20.742 (± 0.741) 18.155 (± 0.670) 22.874 (± 3.780) 20.271 (± 2.590)
MAE_FR_LP 22.589 (± 0.406) 23.406 (± 11.682) 14.296 (± 0.474) 21.585 (± 0.763) 16.781 (± 0.677) 14.440 (± 0.641) 18.582 (± 0.322) 22.147 (± 3.725) 19.228 (± 2.336)
MAE_FR_FT 20.505 (± 0.334) 24.018 (± 13.502) 12.466 (± 0.425) 20.842 (± 0.548) 16.069 (± 0.860) 12.907 (± 0.426) 19.371 (± 0.414) 22.422 (± 4.166) 18.575 (± 2.584)
Table 4: Cross-dataset generalization performance. Models are trained on labeled data from all but five datasets (see Sec 4), and evaluated on held-out datasets to assess OOD generalization. Trait-wise performance includes $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$ metrics. In RTM-AE, cbc is not directly predicted but is derived from cm and cp estimates (cm – cp). We bold and underline best and second best scores respectively.
OOD evaluation.
As shown in Table 4, the fine-tuned MAE (MAE-FR-FT) had the highest performance over all other methods across traits, achieving a slight improvement in $R^{2}$ relative to the supervised baseline ($0.31$ vs. $0.24$), along with the lowest average nRMSE. Since many traits are not associated to single-band features but instead arise from complex interactions across multiple regions of the spectrum, MAE provides a strong prior: it enforces the learning of localized correlations and long-range dependencies within hyperspectral signals, by reconstructing both across adjacent tokens and distant tokens. This prior knowledge facilitates better generalization and more efficient fine-tuning with MAE-FR-FT for the downstream regression. However, this prior alone is not sufficient. When only linear probing is applied (MAE-FR-LP), the model retains general spectral trends leading to underperformance. The necessity of fine-tuning becomes evident in our feature attribution analysis (Fig. 15), where we compared gradient amplitudes across spectral bands for MAE-FR-LP, last block fine-tuning, and full fine-tuning MAE-FR-FT models. While MAE-FR-LP exhibited diffuse and noisy attributions across broad spectral regions, fine-tuning progressively reduced gradient variance, yielding sharper and more interpretable feature importance profiles. This indicates that fine-tuning allows the pretrained prior to be refined toward trait-relevant spectral dependencies, transforming general correlations into targeted representations that drive improved predictive performance.
Other competing methods, such as SR-GAN and RTM-AE, provided modest gains over the supervised baseline. The corresponding scatter plot of the observed and predicted trait values from the different methods is presented in Fig. 14 in the Appendix.
7 Conclusions and perspectives
In this study, we introduce GreenHyperSpectra, a large-scale cross-sensor and cross-ecosystem spectral dataset designed to train machine learning models for plant trait prediction from hyperspectral data. Leveraging GreenHyperSpectra as a pretraining resource, we demonstrated that models using MAE consistently outperformed all other benchmarked methods, including the fully supervised baseline, across a variety of settings. The adaptability of MAE models enables their application to multi-scale remote sensing platforms, including drone, airborne, and satellite imagery, paving the way to investigate how learned spectral features, derived from a heterogeneous spectral dataset, generalize across varying spatial resolutions. MAEs also serve as a strong foundation for advanced transfer learning architectures aimed at improving predictive performance. While we explore default sensor configurations (VNIR+SWIR and VNIR), extending pretrained encoders to other spectral ranges remains open for future work. The MAE architecture shows promise for cross-domain adaptation across heterogeneous sensing modalities through fine-tuning strategies. We contribute towards global pretraining datasets for spectral embeddings while highlighting critical biases affecting generalization. Despite improvements, run-to-run variance reveals challenges in learning stable representations from various ecological data distributions. Future research should expand multi-domain spectral datasets across biomes and sensing conditions to enhance transferability and address geographical and ecosystem-level biases in annotated data. Nevertheless, our pretrained models from GreenHyperSpectra will remain valuable as labeled data from underrepresented regions increases. This study confirms that semi- and self-supervised methods with large-scale pretraining are essential for advancing ecosystem monitoring.
Acknowledgments
We thank all data owners for sharing the data either by request or through the public Ecological Spectral Information System (EcoSIS), Data Publisher for Earth and Environmental Science (PANGEA) and DRYAD platforms. EC, AO and DR acknowledge support for this work from IVADO and the Canada CIFAR AI Chairs program, and computational support from Mila – Quebec AI Institute, including in-kind support from Nvidia Corporation. EC and HF acknowledge the financial support by the Federal Ministry of Education and Research of Germany and by the Sächsische Staatsministerium für Wissenschaft Kultur und Tourismus in the program Center of Excellence for AI-research "Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig", project identification number: ScaDS.AI.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: we summarize four key contributions in Section 1, which are addressed through our experimental design and results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section  7
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This work does not include theoretical analysis.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We share the data and code and provide more details on the experiments in Appendix B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We submit the code and data URL on the OpenReview platform.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Key training details, including model architecture, loss functions, data splits and major modifications, are summarized in the main text. Comprehensive descriptions of training procedures and hyperparameters are provided in Appendix  B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We evaluate all models using three random seeds and report the mean and standard deviation of the performance metrics (i.e. $R^{2}$, nRMSE) to reflect variability due to random initialization. This is stated in the main text and reflected in relevant figures and tables (e.g. Figures 5 and  4). For the cross-dataset OOD evaluation, we report results from a single run due to computational constraints; however, we attempt to account for variability through controlled random sampling, as detailed in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide detailed information on compute resources in Appendix B, including the type of GPUs, model-specific runtime estimates, and the number of trainable parameters per method (see Table 19).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS ethical guidelines and ensured that the paper complies with them.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the positive societal impacts of our work in Section  1 and  7. While we did not identify any clear negative impacts, we remain open to addressing potential concerns if raised during the review process.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The data and models presented in this paper do not pose risks of misuse. All datasets are derived from publicly available, non-sensitive sources, and the models are focused on ecological applications.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets used in this work are publicly available and properly credited in the paper. For each source included in GreenHyperSpectra, we provide citation information, access links in Table 7.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The GreenHyperSpectra dataset introduced in this paper is documented in detail in Appendix A. The appendix includes information about data sources, sensor specifications, geographic and temporal coverage, preprocessing steps, and licensing terms for each included dataset. The asset is composed entirely of publicly available data, contains no personal or sensitive information, and respects the original licenses of the sources.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve any crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methods and experiments presented in this paper do not involve the use of large language models (LLMs).
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Table of figures:
Figure Description
1 Teaser illustration of the proposed semi-/self-supervised frameworks for multi-trait regression.
2 Comparison of GreenHyperSpectra and the labeled dataset, highlighting broader coverage in vegetation types and sensor diversity.
3(a), 3(b) and 3(c) Overview of the SR-GAN, RTM-AE, and MAE architectures for trait prediction.
4 Effect of increasing labeled data volume on R2 performance.
5 Effect of increasing unlabeled pretraining data volume on R2 performance.
6, 7,8 and 9 Dataset characteristics, focusing on spectral variability.
10 and 11 Complementary results to Figs. 4 and  5: nRMSE performance trends with increasing labeled and unlabeled data.
12 and 13 Complementary results to Table 3: Heatmaps of R2 and nRMSE across traits in the half-range input settings to show the performance of MAE vs baseline.
14 Complementary results to Table 4: Observed vs. predicted plots, showing trait-wise calibration in the OOD setting.
15 Feature importance of MAE-based downstream regression as a function of fine-tuning depth (linear probing, final block, and full fine-tuning).
Table 5: Summary of figures and their descriptions.
Table of tables:
Table Description
1 Summary of sensor and platform specifications in GreenHyperSpectra.
2 Trait-wise performance (R2 and nRMSE) of all models under full-range input settings.
3 Trait-wise performance (R2 and nRMSE) of all models under half-range (VNIR) input settings.
4 Trait-wise performance (R2 and nRMSE) of all models under out-of-distribution (OOD) settings.
7 and 8 Dataset details: spectral data characteristics and trait distribution across sources.
9 and 10 Architecture and hyperparameters of SR-GAN.
11, 12 and 13 Architecture, hyperparameters, and RTM configuration of RTM-AE.
14 and 15 Architecture and hyperparameters of MAE.
16, 17 and 18 MAE ablation studies: effects of transformer depth, loss weighting, and token size on trait prediction (R2 and nRMSE).
19 Model size, runtime, and GPU usage across methods.
20, 21, 22, 23, 24 and 25 Complement to Table 4 and Fig. 14: OOD model performance when one vegetation class is excluded from the test set.
26, 27, 28, 29 and 30 Robustness evaluation under additive Gaussian noise during inference, reported for all models (R2 and nRMSE).
Table 6: Summary of tables and their descriptions.
Appendix A Details about the datasets
The data are publicly available here.
Spectral preprocessing.
For standardized cross-instrument comparison, all reflectance spectra were resampled to a uniform wavelength grid spanning the 400-2500 nm solar-reflective range. Spectral measurements were linearly interpolated to an interval of 1 nm, resulting in 2101 bands per sample. Regions of strong atmospheric water absorption, specifically 1351–1430 nm, 1801–2050 nm, and 2451–2500 nm, were removed to minimize noise and signal loss. The remaining bands were smoothed using a Savitzky-Golay filter with a 65 nm window [savitzky1964smoothing]. After these steps, 1721 spectral bands were retained for analysis, providing a high-quality input space for training and evaluation.
Figure 6: Spectral reflectance across wavelengths. This plot shows the variation in canopy reflectance within GreenHyperSpectra across different data sources, highlighting differences due to acquisition conditions and sensor modalities. The colored ranges refer to the visible region.
Figure 7: Sample distribution across vegetation types in GreenHyperSpectra. The plot shows the number of samples in GreenHyperSpectra (blue) and the existing labeled (purple) for each vegetation type, highlighting class imbalance and the relative scarcity of labeled data in certain categories. The vegetation type information was retried from the ESA WORLDCOVER product
Dataset Platform Sensor GSD #Bands Range (nm) Year #Samples Processing Land Cover Source
DB1[Dennison2018RangeCreek] proximal ASD FieldSpec Pro N/A 2151 350–2500 2007 31 Reflectance spectra desert, shrubland Link
2009 49
DB2[DennisonGardner2018Hawaii] proximal ASD FieldSpec FR N/A 1063 352–2476 2000 792 Reflectance spectra forest, shrubland: native-dominated Hawaiian forest types Link
DB3[DennisonRoberts2018SantaMonica] proximal ASD FieldSpec FR N/A 1075 350–2498 1995 226 Reflectance spectra shrubland Link
1996 93
1997 132
1998 10
DB4[Serbin2019UWBNL] proximal SpecEvo PSM3500 N/A 2151 350–2500 2013 6 Reflectance spectra Forest, Gra4Sand, Shrubland, Crops Link
ASD Fieldspec 3 2013 7
ASD Fieldspec 4 2013 49
SpecEvo_PSM3500 2014 60
DB5[Zesati2019ABoVE] proximal SVC HR-1024i 1 m 2178 338–2516 2018 112 Reflectance spectra tundra Link
DB6[Thompson2021EMITReflectance] proximal ASD FieldSpec 3 N/A 2151 350–2500 2001 112 Reflectance spectra urban vegetation Link
DB7[Dennison2019FractionalCover] proximal ASD FieldSpec Pro N/A 210 400–2490 715 Reflectance spectra crops, gra4Sand, forest, shrubland Link
DB8[Herold2004UrbanSB] proximal ASD FieldSpec 3 N/A 1075 350–2498 2001 37 Reflectance spectra urban vegetation Link
DB9[Kokaly2017USGSSpectralLib] proximal ASD FieldSpec 4 Hi-Res NG N/A 2151 350–2500 87 Reflectance spectra Link
DB10[Kokaly2017USGSSpectralLib] proximal Beckman 5270 N/A 480 205–2976 19 Reflectance spectra Link
DB11[Fang2023AquaticVeg] proximal SVC HR-1024i 40 cm 995 346–2499 133 Reflectance spectra aquatic vegetation Link
DB12[Unger2021ArcticMoss] proximal SVC HR-1024i 8 cm 994 338–2515 34 Reflectance spectra tundra Link
DB13[Serbin2015NGEE] proximal SVC HR-1024i N/A 2150 350–2500 2015 44 Reflectance spectra coastal, wetland Link
DB14[Roupioz2023CAMCATT] proximal ASD N/A 2101 400–2500 2021 45 Reflectance spectra urban vegetation Link
DB15[Dennison2018RushValley] proximal ADS FieldSpec Pro N/A 2151 350–2500 2005 82 Reflectance spectra shrubland, steppe Link
DB16[Zesati2019ABoVE] proximal SVC HR-1024i 1 m 994 338–2516 2017 1660 Reflectance spectra tundra Link
DB17[Thompson2021EMITVeg] proximal ASD FieldSpec 3 N/A 2151 350–2500 2001 490 Reflectance spectra coastal, forest, shrubland Link
DB18[Queally2018CAVeg] airborne AVIRIS Classic 17–20 m 244 365–2496 2013 341 ACSR Urban, chaparral, oak woodland, conifer forest Link
2014 37
2016 31
DB19[Kokaly2017USGSSpectralLib] airborne AVIRIS 17–20 m 224 365–2496 32 Link
DB20[PazKagan2015Multiscale] airborne AisaFenix 1 m 360 400–2400 2014 22889 ACSR Forest, Ecology, Land Cover, Agriculture Link
DB21[Campbell2016Mongu] spaceborne Hyperion 30 m 220 400–2500 2008 2009 25 ACSR forest Link
DB22[Weinstein2022NeonTree] airborne NEON AOP 1 m 426 380–2510 2018 10322 ACSR Link
DB23[hu2023mdas] airborne Hyspex 368 417–2484 2018 9993 ACSR crops Link
DB24[Weinstein2022NeonTree] airborne NEON AOP 1 m 426 380–2510 2019 16373 ACSR Link
DB25[Vivone2022Fusion] spaceborne PRISMA 30 m 69 400–2500 2021 2155 ACSR Link
DB26[Vivone2022Fusion] spaceborne PRISMA 30 m 63 400–2500 2021 10000 ACSR
DB27[fuchs2023hyspecnet] spaceborne EnMAP 30 m 224 418–2445 2022 1890 ACSR Link
DB28 airborne NEON AOP 1 m 426 380–2510 2022 3959 ACSR Link
DB29[Roupioz2023CAMCATT] airborne AisaFenix 1 m 420 382–2499 2021 31811 ACSR urban vegetation Link
DB30[Chadwick2025SHIFT] airborne Aviris NG 20 m 425 380–2510 2022 911 ACSR Mediterranean ecosystem Link
DB31 spaceborne EMIT 60 m 285 381–2492 2024 410 ACSR Link [^2] Contains modified EMIT data, the original data used is licensed under the Apache License, Version 2.0.
DB32 spaceborne EnMAP 30 m 224 418–2445 2022- 2024 6653 ACSR temperate forest EnMAP [^1] Contains modified EnMAP data © DLR [2024]. See EnMAP Portal.
DB33 spaceborne EnMAP 30 m 224 418–2445 2022- 2024 1655 ACSR Mediterranean ecosystem EnMAP [^1] Contains modified EnMAP data © DLR [2024]. See EnMAP Portal.
DB34 spaceborne EnMAP 30 m 224 418–2445 2022- 2024 2088 ACSR temperate gra4Sand EnMAP [^1] Contains modified EnMAP data © DLR [2024]. See EnMAP Portal.
DB35 spaceborne EnMAP 30 m 224 418–2445 2022- 2024 4846 ACSR tropical forest EnMAP [^1] Contains modified EnMAP data © DLR [2024]. See EnMAP Portal.
DB36 spaceborne EnMAP 30 m 224 418–2445 2022- 2024 6337 ACSR tropical savanna EnMAP  [^1] Contains modified EnMAP data © DLR [2024]. See EnMAP Portal.
Table 7: Summary of data sources of GreenHyperSpectra. Technical detailed on the collected spectra and their corresponding sources. GSD = Ground Sampling Distance. ACSR = Atmospherically corrected surface reflectance.
Figure 8: t-SNE projection of reflectance spectra. Each subplot shows the projected spectral signatures from a specific data source in GreenHyperSpectra, illustrating variability driven by differences in sensors, biomes, or acquisition conditions. For each source, spectra from GreenHyperSpectra are compared to those in the aggregated annotated dataset. Orange points represent labeled spectra, and blue points denote unlabeled samples.
Figure 9: Spatial distribution of a subset from GreenHyperSpectra vs annotated data. Points represent sample locations of the existing annotated data (left) and the GreenHyperSpectra subset (right). This subset, comprising 80,000 samples, was selected from the full dataset to ensure broad coverage of geographic regions and acquisition conditions. It is used for sample sensitivity analysis to assess the impact of data quantity on model performance, while enabling computationally efficient experimentation.
Table 8: Descriptive statistics for plant traits in the aggregated annotated dataset[cherif2023spectra]. List of traits and units: leaf mass per area ($\text{g}/\text{cm}^{2}$) = Cm, leaf protein content ($\text{g}/\text{cm}^{2}$) = Cp, equivalent water thickness (cm) = Cw, leaf total chlorophyll content ($\mu\text{g}/\text{cm}^{2}$) = Cab, leaf carotenoid content ($\mu\text{g}/\text{cm}^{2}$)= Car, leaf anthocyanin content ($\mu\text{g}/\text{cm}^{2}$) = Anth, Leaf Area Index ($\text{m}^{2}/\text{m}^{2}$) = LAI and carbon-based constituents ($\text{g}/\text{cm}^{2}$)= cbc (Cm-Cp).
Trait Count Mean Std Min 25% 50% 75% Max
cab 2593 39.1234 14.2312 4.4483 28.2500 38.0042 49.0675 229.4975
cw 2782 0.0160 0.0166 0.0000 0.0096 0.0130 0.0184 0.5138
cm 4062 0.0101 0.0083 0.0000 0.0051 0.0080 0.0117 0.0682
LAI 1656 3.4927 1.7178 0.0633 2.1944 3.4691 4.7743 8.7700
cp 3031 0.0009 0.0005 0.0000 0.0006 0.0008 0.0011 0.0050
cbc 3031 0.0104 0.0086 0.0000 0.0056 0.0078 0.0123 0.0671
car 1873 8.6925 2.8232 1.1826 6.9679 8.5176 10.2998 40.4432
anth 644 1.2730 0.4095 0.5610 0.9491 1.2345 1.5226 2.9811
Appendix B Details about Models
The code for accessing the dataset and benchmarking experiments can be found here. The trained model objects are also available here
B.1 Semi-supervised regression generative adversarial network (SR-GAN)
To enable trait prediction from unlabeled hyperspectral spectra, we adopt a semi-supervised GAN framework [olmschenk2019generalizing]. The generator $G$ learns to synthesize spectra that are indistinguishable from real data by optimizing $\mathcal{L}_{\text{gen}}$, while the discriminator $D$ learns both to distinguish real from fake spectra and to regress plant trait values from real labeled data optimizing $\mathcal{L}_{disc}$ . We adopt this notation for $D$ outputs: $f$ is an intermediate feature extractor and $D\circ f$ is final layer trait prediction. $\text{Dist}(\cdot,\cdot)$ denotes a distance metric (e.g., cosine or Euclidean).
Notations: $x_{\text{fake}}$: generated fake spectra from the generator; $x_{\text{unlb}}$: unlabeled sample from GreenHyperSpectra; $x_{\text{lb}}$: spectra sample from the labeled data;
Generator Matching Loss.
The generator is trained to align the generated spectra with real spectra in the latent feature space:
$\mathcal{L}_{\text{gen}}=\lambda_{\text{gen}}\cdot\text{Dist}\left(f(x_{\text{fake}}),f(x_{\text{unlabeled}})\right),$ (1)
where $\lambda_{\text{gen}}$ controls the influence of the generator loss.
Labeled Supervised Loss.
For labeled spectra $x_{\text{labeled}}$ with corresponding trait references $y$, we define a standard supervised regression loss:
$\mathcal{L}_{\text{labeled}}=\lambda_{\text{labeled}}\cdot\text{MSE}(D\circ f(x_{\text{labeled}}),y),$ (2)
where $\lambda_{\text{labeled}}$ is a weighting coefficient, and MSE denotes the mean squared error between predicted and true traits.
Unlabeled Matching Loss.
To regularize the feature space, we encourage the feature representations $f(\cdot)$ extracted by $D$ from labeled and unlabeled real spectra to be similar:
$\mathcal{L}_{\text{unlabeled}}=\lambda_{\text{unlabeled}}\cdot\lambda_{\text{srgan}}\cdot\text{Dist}\left(f(x_{\text{labeled}}),f(x_{\text{unlabeled}})\right),$ (3)
where and $\lambda_{\text{unlabeled}}$ and $\lambda_{\text{srgan}}$ are scaling factors.
Fake Contrastive Loss.
The discriminator is further trained to push away fake spectra $x_{\text{fake}}=G(z)$ from real ones in the feature space:
$\mathcal{L}_{\text{fake}}=\lambda_{\text{fake}}\cdot\lambda_{\text{srgan}}\cdot\text{Dist}\left(f(x_{\text{unlabeled}}),f(x_{\text{fake}})\right),$ (4)
where $\lambda_{\text{fake}}$ weights the contrastive term.
Gradient Penalty.
A gradient penalty is used to enforce Lipschitz continuity, stabilizing the training of the discriminator:
$\mathcal{L}_{\text{GP}}=\lambda_{\text{GP}}\cdot\mathbb{E}_{\hat{x}}\left[max(0,\left(\left\|\nabla_{\hat{x}}f(\hat{x})\right\|_{2}-1\right)^{2})\right],$ (5)
where $\hat{x}$ is an interpolated sample between $x_{\text{fake}}$ and $x_{\text{unlabeled}}$, and $\lambda_{\text{GP}}$ is the penalty coefficient.
Total Losses.
The complete losses for the discriminator are defined as:
$\mathcal{L}_{disc}=\mathcal{L}_{\text{labeled}}+\mathcal{L}_{\text{unlabeled}}+\mathcal{L}_{\text{fake}}+\mathcal{L}_{\text{GP}},$ (6)
Table 9: Architectural details of the convolutional GAN model used for spectral generation and trait regression.
Network Layer Description
Generator Input Latent vector $z\in\mathbb{R}^{d}$
Fully Connected Linear: $d\rightarrow 64\times\frac{S}{4}$, reshaped to $(64,\frac{S}{4})$
Transposed Conv1D $(64,\frac{S}{4})\rightarrow(64,S)$, kernel=16, stride=4, pad=6
Residual Stack Three residual blocks (dilations=1, 3, 9), LeakyReLU, skip connections
Output Conv1D Conv1D: $(64,S)\rightarrow(1,S)$, kernel=7, pad=3
Output Activation Tanh activation to constrain to [1,1][-1,1]
Discriminator Input Spectral input $x\in\mathbb{R}^{1\times S}$
Conv1D Layer 1 SpectralNorm: $(1,S)\rightarrow(128,S/2)$, kernel=3, stride=2, pad=1
BatchNorm + Activation BatchNorm1D + LeakyReLU
Conv1D Layer 2 SpectralNorm: $(128,S/2)\rightarrow(128,S/4)$
Conv1D Layer 3 (output) SpectralNorm: $(128,S/4)\rightarrow(128,S/8)$
Adaptive Pooling AdaptiveAvgPool1D (optional)
Flatten Flatten to $(128\times S/8)$
Dropout Dropout $p=0.4$
Fully Connected (output) Linear: $128\times S/8\rightarrow n_{\text{traits}}$
Table 10: Training hyperparameters for the SR-GAN. This table lists the default hyperparameters and optimization settings used during training for both generator and discriminator components.
Parameter Value Description
input shape 1720 or 500 Number of spectral input bands
latent dim 100 Generator latent vector size
n_lb 8 Number of predicted plant traits
batch size 128 Samples per batch
n_epochs 300 Total training epochs
learning rate G 1e-4 Generator optimizer learning rate
learning rate D 4*1e-4 Discriminator optimizer learning rate
optimizers Adam (amsgrad=True) Optimizer
weight decay 1e-4 L2 regularization
lambda_fk 1.0 Generator adversarial loss weight
lambda_un 10.0 Unsupervised feature loss weight
labeled_loss_multiplier 1.0 Supervised regression loss weight
matching_loss_multiplier 1.0 Real/fake match loss weight
contrasting_loss_multiplier 1.0 Contrastive loss weight
srgan_loss_multiplier 1.0 Contrastive loss weight
gradient penalty on True Enable gradient penalty
gradient_penalty_multiplier 10.0 Weight for GP term
augmentation True Data augmentation
contrasting_distance_function CosineEmbeddingLoss Real/fake separation
matching_distance_function CosineEmbeddingLoss Real-real alignment
labeled_loss_function Huber loss Regression loss for traits
B.2 Model Architecture Description for RTM-AE
Table 11: Architecture of the RTM-AE model.
Module Layer Description
Encoder Input Spectral input $x\in\mathbb{R}^{1\times S}$
Fully Connected 1 $S\rightarrow 64$, followed by LayerNorm and ReLU
Fully Connected 2 $64\rightarrow 32$, followed by LayerNorm and ReLU
Fully Connected 3 $32\rightarrow 16$, followed by LayerNorm and ReLU
Trait Output Layer $16\rightarrow n_{\text{traits}}$
RTM Decoder Non-learnable Module PROSAIL-PRO: $n_{\text{traits}}\rightarrow\tilde{x}\in\mathbb{R}^{1\times 2101}$
Correction Block Fully Connected 1 $2101\rightarrow 8404$, followed by ReLU
Fully Connected 2 $8404\rightarrow\hat{x}\in\mathbb{R}^{1\times 2101}$
Output Corrected reflectance spectrum in $\mathbb{R}^{S}$
Table 12: Training hyperparameters for the RTM-AE. This table lists the default hyperparameters and optimization settings used during training.
Parameter Value Description
input shape 1720 or 500 Number of spectral input bands
latent dimension 8 Number of biophysical traits (latent features)
output spectrum length 2101 Number of bands in RTM-simulated output
batch size 128 Number of samples per training batch
training epochs 300 Set during experimental runs
learning rate 1e-4 Learning rate used for the Adam optimizer
weight decay 1e-4 L2 regularization term
optimizer Adam (amsgrad=True) Optimizer used
reconstruction loss Cosine similarity + MAE Match predicted vs. input spectra
label loss Huber loss Trait prediction loss on labeled samples
gradient stabilization Enabled Replace gradients when $\|\nabla\|<10^{-5}$
RTM decoder PROSAIL-PRO Fixed physics-based decoder
leaf model PROSPECTPro Leaf optical model
canopy model SAIL Canopy RTM model
Table 13: Parameter configuration for the PROSAIL-PRO model [feret2021prospect]. This table presents the default settings and corresponding notations for parameters used in the RTM block, which simulates leaf and canopy reflectance based on the PROSPECT-PRO [jacquemoud1990prospect] and 4SAIL [verhoef2007unified] models.
Model Variable Notation (unit) Range
[…]
Water content Cw (g/m2) Variable
Protein content Cp (g/m2) Variable
Carbon-based constituents CBC (g/m2) = Cm - Cp Variable
Brown pigment content Brown (–) 0.25
Structural coefficient Ns (–) 1.5
4SAIL Leaf area index LAI (m2/m2) Variable
Average leaf inclination angle LIDF (Beta index) 5
Fraction of dry soil psoil 0.8
Hotspot hspot 0.01
Viewing zenith angle tto () 0
Solar zenith angle $t$s () 30
Relative azimuth angle psi () 0
B.3 Model Architecture Description for the 1D masked autoencoder (MAE)
B.3.1 Training description
Table 14: Architecture of the Masked Autoencoder (MAE)
Module Layer Description
Input Input Spectrum 1D spectral vector $x\in\mathbb{R}^{1\times S}$
Patch Embedding Tokens of size $T$ via frozen Conv1D, $N=S/T$
Positional Embedding Fixed 1D sin-cos embeddings
Masking Random masking of 75% of tokens
Encoder Transformer Blocks $d$ blocks with $h$-head attention, MLP ($4\times d$)
Normalization LayerNorm
Output Latent representation $\in\mathbb{R}^{N\times d}$
Decoder Linear Projection Projects latent dim to decoder dimensions
Token Restoration Restore with mask tokens using ids_restoreids\_restore
Transformer Blocks $d^{\prime}$ blocks with $h^{\prime}$-head attention
Output Projection Linear: decoder dim $\rightarrow$ token size $T$
Reconstruction Spectrum Output Reconstructed full spectrum $\in\mathbb{R}^{1\times S}$
Table 15: Training hyperparameters of the MAE. This table lists the default hyperparameters and optimization settings used during training on the pretext task (spectra reconstruction).
Parameter Value Description
Input dimension ($S$) 1720 or 500 Number of spectral bands
Patch size ($T$) 10, 20, 40, 430 Token size (Ablation study)
Embedding dimension ($d$) 128 Latent dimension
Encoder depth ($d$) 4, 6, 8 Transformer blocks (Ablation study)
Encoder heads ($h$) 4, 8, 16 Attention heads (Ablation study)
Decoder depth ($d^{\prime}$) 4 Decoder transformer depth
Decoder heads ($h^{\prime}$) 4 Decoder attention heads
Mask ratio 0.75 Fraction of masked tokens
Loss function WLoss * Cosine + MSE Shape and amplitude combined loss
WLoss 1, 0.1, 0.001, 0 Weight for cosine term
MLP ratio 4.0 MLP expansion factor
Attention dropout 0.0 Dropout in attention
Projection dropout 0.0 Dropout in projections
Optimizer AdamW With AMSGrad
Learning rate 5e-4 Initial learning rate
Weight decay 1e-4 L2 regularization
Batch size 128 Samples per batch
Epochs 500 Total training epochs
B.3.2 Ablation study results
Table 16: Ablation study on the effect of transformer depth and attention heads in the MAE model. This table reports the $R^{2}$ and nRMSE values for MAE models evaluated on the trait prediction task, with varying transformer depths and numbers of attention heads. The highest $R^{2}$ scores and lowest nRMSE values are highlighted.
Depth Heads Final Val $R^{2}$ Final Val nRMSE
6 4 0.4492 15.73
6 8 0.4090 16.31
6 16 0.4327 15.96
8 4 0.4351 15.90
8 8 0.4356 15.90
8 16 0.3937 16.44
10 4 0.4060 16.30
10 8 0.4092 16.28
10 16 0.4692 15.43
Table 17: Ablation study on the effect of cosine similarity loss weight in the MAE model. This table presents the $R^{2}$ and nRMSE values for trait prediction as the cosine similarity loss weight ($w_{\text{loss}}$) is varied in the MAE objective. The highest $R^{2}$ scores and lowest nRMSE values are highlighted.
$w_{\text{loss}}$ Final Val $R^{2}$ Final Val nRMSE
1 0.5018 14.96
0.1 0.4907 15.15
0.01 0.4233 16.08
0.001 0.4627 15.55
0 0.4698 15.42
Table 18: Ablation study on the effect of spectral token size in the MAE model. This table reports the $R^{2}$ and nRMSE values for trait prediction as the spectral token (sequence) size is varied in the MAE architecture. The highest $R^{2}$ scores and lowest nRMSE values are highlighted.
Token Size Final Val $R^{2}$ Final Val nRMSE
10 0.4542 15.68
20 0.5018 14.96
40 0.4744 15.35
430 0.2683 18.05
Appendix C Resource requirements
Table 19: Comparison of methods by model size, runtime, and hardware usage. This table summarizes the number of trainable parameters, average runtime, and GPU type used for each method, providing insights into their computational efficiency and resource requirements.
Method # Trainable Parameters Run Time GPU
Supervised (EffNetB0) 6,998,280 $\sim$11h Quadro RTX 8000
MAE (pretext task) Encoder: 2,006,912 $\sim$20h Quadro RTX 8000
Decoder: 1,220,116
Total: 3,227,028
MAE (downstream Linear Probing) 1,288 $\sim$15 min Quadro RTX 8000
MAE (downstream Fine Tuning) 1,607,196 $\sim$15 min Quadro RTX 8000
SR-GAN Discriminator: 319,496 $\sim$2.5d Quadro RTX 8000
Generator: 2,920,130
Total: 3,239,626
RTM-AE 35,437,289 $\sim$16h NVIDIA L40S
Appendix D complements to results
Figure 10: Evaluation of trait prediction with variable-size labeled sets. Validation performance (nRMSE) as a function of the percentage of labeled data used for training. The average nRMSE performance across all traits is indicated by the dashed box. Lower nRMSE values indicate better predictive performance.
Figure 11: Evaluation of trait prediction with variable-size unlabeled sets. Validation performance (nRMSE) as a function of the percentage of labeled data used for training. The average nRMSE performance across all traits is indicated by the dashed box. Lower nRMSE values indicate better predictive performance.
Figure 12: Trait-wise performance heatmaps in the half-range (HR) for MAE vs Baseline. The heatmap displays the coefficient of determination ($R^{2}$; higher is better). Each cell represents the average performance across runs for a given trait-method combination.
Figure 13: Trait-wise performance heatmaps in the half-range (HR) setting MAE vs Baseline. The heatmap displays the normalized root mean square error (nRMSE; lower is better). Each cell represents the average performance across runs for a given trait-method combination.
Figure 14: Observed vs. predicted trait values in the cross-dataset OOD setup. Each subplot corresponds to a specific trait (rows) and method (columns), comparing predicted values to reference data. The black line indicates the 1:1 reference. $R^{2}$ and nRMSE values are reported in each plot to quantify predictive performance.
cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
Supervised 0.2836 0.1107 0.3728 0.1176 0.0795 0.3339 0.1784 0.0810 0.1947
SR_GAN 0.3108 0.3065 0.4931 -0.1324 0.1924 0.4810 0.0831 0.1022 0.2296
RTM_AE 0.1846 0.0577 0.4681 0.0790 0.0704 -0.1646 0.1902 -0.0186 0.1083
MAE_FR_LP 0.0542 0.1803 0.3857 0.1673 0.0100 0.4034 0.1864 0.0442 0.1789
MAR_FR_FT 0.2257 0.1337 0.5184 0.2239 0.2425 0.4837 0.2001 0.0967 0.2656
nRMSE $(\downarrow)$
Supervised 20.1711 30.4925 16.2647 23.4458 18.6372 17.3007 17.7155 20.3835 20.5514
SR_GAN 19.7855 26.9263 14.6236 26.5084 17.4572 15.2720 18.7152 20.1473 19.9294
RTM_AE 21.5192 31.3877 14.9780 23.9535 18.7291 22.8767 17.5886 21.4592 21.5615
MAE_FR_LP 23.1762 29.2746 16.0959 22.7759 19.3277 16.3739 17.6299 20.7873 20.6802
MAR_FR_FT 20.9691 30.0953 14.2528 21.9882 16.9061 15.2323 17.4806 20.2090 19.6417
Table 20: Cross-dataset generalization by vegetation type: Tundra. Unlike the overall OOD results (Table 4), here we exclude samples from tundra during evaluation to assess its individual impact on model generalization. Trait-wise performance is reported using $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$. We highlight the best and second-best scores in bold and underline, respectively.
cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
Supervised 0.3586 0.0432 0.1471 0.1229 -0.0927 0.1366 0.3344 0.0810 0.1414
SR_GAN 0.3645 0.2524 0.2372 -0.1402 0.0366 0.2824 -0.0012 0.1022 0.1417
RTM_AE 0.1744 -0.0311 0.1994 0.1343 -0.0779 -0.3144 0.2837 -0.0186 0.0437
MAE_FR_LP 0.0412 0.1122 0.1686 0.1624 -0.2675 0.2253 0.2823 0.0442 0.0961
MAE_FR_FT 0.2188 0.0491 0.3057 0.2248 0.0710 0.2927 0.3348 0.0967 0.1992
nRMSE $(\downarrow)$
Supervised 18.8612 37.3976 18.4678 23.8133 22.9505 19.2217 17.0518 20.3835 22.2684
SR_GAN 18.7736 33.0584 17.4651 27.1220 21.5493 17.5235 20.9137 20.1473 22.0691
RTM_AE 21.3980 38.8223 17.8924 23.6594 22.7941 23.7159 17.6905 21.4592 23.4290
MAE_FR_LP 23.0604 36.0248 18.2334 23.2713 24.7179 18.2073 17.7073 20.7873 22.7512
MAE_FR_FT 20.8151 37.2836 16.6626 22.3877 21.1619 17.3969 17.0472 20.2090 21.6205
Table 21: Cross-dataset generalization by vegetation type: Forest. Unlike the overall OOD results (Table 4), here we exclude samples from forest during evaluation to assess its individual impact on model generalization. Trait-wise performance is reported using $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$. We highlight the best and second-best scores in bold and underline, respectively.
cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
Supervised 0.2522 0.1019 0.3391 -0.0108 -0.1300 0.2478 0.1712 0.0810 0.1316
SR_GAN 0.2811 0.3282 0.4731 -0.0505 -0.0075 0.3965 0.0441 0.1022 0.1959
RTM_AE 0.2239 0.0859 0.4679 -0.1631 -0.1097 -0.2146 0.1450 -0.0186 0.0521
MAE_FR_LP 0.0786 0.2257 0.3792 -0.0207 -0.1404 0.3455 0.1641 0.0442 0.1345
MAR_FR_FT 0.2897 0.1377 0.5074 0.0678 0.0769 0.4326 0.1835 0.0967 0.2240
nRMSE $(\downarrow)$
Supervised 20.3244 34.6730 16.6769 24.8992 20.4582 18.4717 17.4666 20.3835 21.6692
SR_GAN 19.9301 29.9879 14.8919 25.3229 19.3170 16.5456 18.7583 20.1473 20.6126
RTM_AE 20.7059 34.9801 14.9635 26.7101 20.2731 23.4729 17.7401 21.4592 22.5381
MAE_FR_LP 22.5613 32.1942 16.1629 25.0216 20.5521 17.2313 17.5413 20.7873 21.5065
MAR_FR_FT 19.8086 33.9740 14.3981 23.9122 18.4904 16.0433 17.3367 20.2090 20.5215
Table 22: Cross-dataset generalization by vegetation type: Crops. Unlike the overall OOD results (Table 4), here we exclude samples from crops during evaluation to assess its individual impact on model generalization. Trait-wise performance is reported using $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$. We highlight the best and second-best scores in bold and underline, respectively.
cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
Supervised 0.3099 0.1359 0.4383 0.1063 0.1549 0.4156 0.1811 0.0810 0.2279
SR_GAN 0.3318 0.3229 0.5161 -0.1941 0.2496 0.5148 0.0826 0.1022 0.2408
RTM_AE 0.2156 0.0706 0.5020 0.0455 0.1366 -0.1417 0.1872 -0.0186 0.1247
MAE_FR_LP 0.0860 0.2016 0.4433 0.1270 0.0973 0.4713 0.1868 0.0442 0.2072
MAR_FR_FT 0.2459 0.1419 0.5690 0.2074 0.3095 0.5471 0.2068 0.0967 0.2905
nRMSE $(\downarrow)$
Supervised 20.0381 31.1490 14.8665 23.4980 17.7968 15.6488 17.7263 20.3835 20.1384
SR_GAN 19.7169 27.5735 13.7989 27.0932 16.7698 14.2586 18.7612 20.1473 19.7649
RTM_AE 21.3638 32.3043 13.9977 24.2834 17.9889 21.8728 17.6592 21.4592 21.3661
MAE_FR_LP 23.0615 29.9417 14.7998 23.2242 18.3933 14.8854 17.6639 20.7873 20.3446
MAR_FR_FT 20.9468 31.0410 13.0228 22.1287 16.0873 13.7758 17.4458 20.2090 19.3322
Table 23: Cross-dataset generalization by vegetation type: Shrubland. Unlike the overall OOD results (Table 4), here we exclude samples from shrubland during evaluation to assess its individual impact on model generalization. Trait-wise performance is reported using $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$. We highlight the best and second-best scores in bold and underline, respectively.
cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
Supervised 0.1565 0.2080 0.3163 0.2771 0.0615 0.3349 0.1825 NaN 0.2195
SR_GAN 0.1989 0.3005 0.4212 -0.1232 0.1658 0.4556 0.1364 NaN 0.2222
RTM_AE 0.0098 0.0615 0.4207 0.3146 0.0452 -0.1678 0.2005 NaN 0.1264
MAE_FR_LP -0.1716 0.1244 0.3338 0.4612 -0.0172 0.3976 0.2543 NaN 0.1975
MAR_FR_FT 0.0393 0.2813 0.4818 0.4239 0.2123 0.4890 0.1801 NaN 0.3011
nRMSE $(\downarrow)$
Supervised 21.9248 19.3998 16.7762 20.4917 18.6994 17.1365 20.1887 NaN 19.2310
SR_GAN 21.3651 18.2311 15.4364 25.5063 17.6296 15.5046 20.7497 NaN 19.2033
RTM_AE 23.7549 21.1173 15.4424 19.9577 18.8617 22.7073 19.9653 NaN 20.2581
MAE_FR_LP 25.8392 20.3969 16.5597 17.6901 19.4682 16.3092 19.2815 NaN 19.3636
MAR_FR_FT 23.3975 18.4795 14.6050 18.2936 17.1319 15.0207 20.2180 NaN 18.1638
Table 24: Cross-dataset generalization by vegetation type: Grassland. Unlike the overall OOD results (Table 4), here we exclude samples from grassland during evaluation to assess its individual impact on model generalization. Trait-wise performance is reported using $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$. We highlight the best and second-best scores in bold and underline, respectively.
cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
Supervised 0.2861 0.1110 0.3663 0.1023 0.0663 0.3336 0.1836 0.0810 0.1913
SR_GAN 0.3143 0.3069 0.4736 -0.1479 0.1707 0.4667 0.0895 0.1022 0.2220
RTM_AE 0.1849 0.0567 0.4620 0.0720 0.0592 -0.1583 0.1924 -0.0186 0.1063
MAE_FR_LP 0.0545 0.1804 0.3809 0.1577 -0.0022 0.4005 0.1858 0.0442 0.1752
MAR_FR_FT 0.2283 0.1339 0.5162 0.2179 0.2266 0.4849 0.2077 0.0967 0.2640
nRMSE $(\downarrow)$
Supervised 20.1602 30.5654 16.0148 23.5725 18.9082 16.9249 17.6727 20.3835 20.5253
SR_GAN 19.7600 26.9896 14.5977 26.6337 17.8197 15.1402 18.6633 20.1473 19.9689
RTM_AE 21.5421 31.4855 14.7563 23.9671 18.9805 22.3141 17.5763 21.4592 21.5101
MAE_FR_LP 23.2023 29.3485 15.8300 22.8332 19.5902 16.0532 17.6486 20.7873 20.6617
MAR_FR_FT 20.9645 30.1712 13.9845 22.0142 17.2109 14.8817 17.4092 20.2090 19.6056
Table 25: Cross-dataset generalization by vegetation type: Mix. Unlike the overall OOD results (Table 4), here we exclude samples from mix during evaluation to assess its individual impact on model generalization. Trait-wise performance is reported using $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$. We highlight the best and second-best scores in bold and underline, respectively.
Noise cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
0.01 0.552 $\pm$ 0.010 0.602 $\pm$ 0.021 0.677 $\pm$ 0.006 0.564 $\pm$ 0.030 0.659 $\pm$ 0.008 0.679 $\pm$ 0.008 0.588 $\pm$ 0.006 0.422 $\pm$ 0.062 0.593 $\pm$ 0.019
0.03 0.392 $\pm$ 0.072 0.387 $\pm$ 0.095 0.426 $\pm$ 0.059 0.285 $\pm$ 0.094 0.404 $\pm$ 0.118 0.448 $\pm$ 0.065 0.374 $\pm$ 0.101 0.298 $\pm$ 0.032 0.377 $\pm$ 0.079
0.05 -0.031 $\pm$ 0.214 -0.123 $\pm$ 0.212 -0.011 $\pm$ 0.060 -0.181 $\pm$ 0.230 -0.083 $\pm$ 0.173 0.009 $\pm$ 0.059 -0.127 $\pm$ 0.220 0.028 $\pm$ 0.045 -0.065 $\pm$ 0.152
nRMSE $(\downarrow)$
0.01 16.704 $\pm$ 0.181 13.409 $\pm$ 0.354 10.534 $\pm$ 0.100 17.332 $\pm$ 0.620 10.165 $\pm$ 0.121 10.674 $\pm$ 0.127 12.975 $\pm$ 0.093 16.935 $\pm$ 0.906 13.591 $\pm$ 0.313
0.03 19.436 $\pm$ 1.164 16.617 $\pm$ 1.326 14.021 $\pm$ 0.727 22.154 $\pm$ 1.436 13.392 $\pm$ 1.361 13.983 $\pm$ 0.834 15.952 $\pm$ 1.323 18.669 $\pm$ 0.422 16.778 $\pm$ 1.074
0.05 25.244 $\pm$ 2.650 22.482 $\pm$ 2.104 18.624 $\pm$ 0.555 28.434 $\pm$ 2.725 18.078 $\pm$ 1.438 18.753 $\pm$ 0.563 21.385 $\pm$ 2.152 21.969 $\pm$ 0.507 21.871 $\pm$ 1.587
Table 26: Supervised: Noise robustness analysis. Model performance under different noise intensities (0.01, 0.03, 0.05). Trait-wise $R^{2}$ (higher is better) and nRMSE (lower is better) are reported as mean $\pm$ standard deviation.
Noise cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
0.01 0.539 $\pm$ 0.018 0.544 $\pm$ 0.006 0.630 $\pm$ 0.010 0.522 $\pm$ 0.012 0.528 $\pm$ 0.017 0.659 $\pm$ 0.007 0.533 $\pm$ 0.009 0.497 $\pm$ 0.012 0.556 $\pm$ 0.011
0.03 0.329 $\pm$ 0.047 0.292 $\pm$ 0.031 0.318 $\pm$ 0.060 0.375 $\pm$ 0.025 0.264 $\pm$ 0.047 0.348 $\pm$ 0.096 0.283 $\pm$ 0.054 0.108 $\pm$ 0.118 0.290 $\pm$ 0.060
0.05 -0.038 $\pm$ 0.164 -0.071 $\pm$ 0.122 -0.195 $\pm$ 0.300 0.065 $\pm$ 0.212 -0.246 $\pm$ 0.280 -0.196 $\pm$ 0.364 -0.080 $\pm$ 0.103 -0.468 $\pm$ 0.317 -0.154 $\pm$ 0.233
nRMSE $(\downarrow)$
0.01 16.943 $\pm$ 0.323 14.367 $\pm$ 0.097 11.272 $\pm$ 0.147 18.143 $\pm$ 0.236 11.958 $\pm$ 0.218 11.008 $\pm$ 0.115 13.810 $\pm$ 0.133 15.811 $\pm$ 0.185 14.164 $\pm$ 0.182
0.03 20.421 $\pm$ 0.708 17.887 $\pm$ 0.390 15.294 $\pm$ 0.664 20.773 $\pm$ 0.442 14.897 $\pm$ 0.478 15.175 $\pm$ 1.130 17.108 $\pm$ 0.656 21.016 $\pm$ 1.414 17.821 $\pm$ 0.735
0.05 25.367 $\pm$ 1.969 21.970 $\pm$ 1.256 20.147 $\pm$ 2.468 25.295 $\pm$ 2.775 19.220 $\pm$ 2.021 20.449 $\pm$ 3.018 20.982 $\pm$ 1.001 26.895 $\pm$ 3.012 22.541 $\pm$ 2.190
Table 27: SR_GAN: Noise robustness analysis. Model performance under different noise intensities (0.01, 0.03, 0.05). Trait-wise $R^{2}$ (higher is better) and nRMSE (lower is better) are reported as mean $\pm$ standard deviation.
Noise cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
0.01 0.576 $\pm$ 0.028 0.633 $\pm$ 0.020 0.659 $\pm$ 0.018 0.549 $\pm$ 0.029 0.654 $\pm$ 0.012 0.666 $\pm$ 0.022 0.551 $\pm$ 0.040 0.312 $\pm$ 0.140 0.575 $\pm$ 0.038
0.03 0.527 $\pm$ 0.037 0.595 $\pm$ 0.004 0.620 $\pm$ 0.032 0.515 $\pm$ 0.043 0.602 $\pm$ 0.015 0.636 $\pm$ 0.029 0.499 $\pm$ 0.056 0.219 $\pm$ 0.164 0.527 $\pm$ 0.048
0.05 0.377 $\pm$ 0.023 0.439 $\pm$ 0.082 0.456 $\pm$ 0.054 0.457 $\pm$ 0.042 0.456 $\pm$ 0.068 0.456 $\pm$ 0.066 0.350 $\pm$ 0.055 -0.258 $\pm$ 0.880 0.342 $\pm$ 0.159
nRMSE $(\downarrow)$
0.01 16.245 $\pm$ 0.531 12.892 $\pm$ 0.346 10.817 $\pm$ 0.281 17.615 $\pm$ 0.567 10.246 $\pm$ 0.172 10.879 $\pm$ 0.350 13.536 $\pm$ 0.616 18.420 $\pm$ 1.940 13.831 $\pm$ 0.600
0.03 17.153 $\pm$ 0.667 13.543 $\pm$ 0.066 11.411 $\pm$ 0.486 18.302 $\pm$ 0.782 10.987 $\pm$ 0.214 11.356 $\pm$ 0.451 14.282 $\pm$ 0.813 19.619 $\pm$ 2.134 14.582 $\pm$ 0.702
0.05 19.690 $\pm$ 0.362 15.899 $\pm$ 1.198 13.647 $\pm$ 0.686 19.385 $\pm$ 0.771 12.828 $\pm$ 0.818 13.884 $\pm$ 0.863 16.284 $\pm$ 0.692 24.058 $\pm$ 8.329 16.959 $\pm$ 1.715
Table 28: RTM_AE: Noise robustness analysis. Model performance under different noise intensities (0.01, 0.03, 0.05). Trait-wise $R^{2}$ (higher is better) and nRMSE (lower is better) are reported as mean $\pm$ standard deviation.
Noise cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
0.01 0.518 $\pm$ 0.003 0.432 $\pm$ 0.005 0.587 $\pm$ 0.005 0.416 $\pm$ 0.006 0.438 $\pm$ 0.009 0.591 $\pm$ 0.006 0.443 $\pm$ 0.008 0.170 $\pm$ 0.017 0.449 $\pm$ 0.007
0.03 0.153 $\pm$ 0.029 0.357 $\pm$ 0.026 0.486 $\pm$ 0.004 0.395 $\pm$ 0.011 0.253 $\pm$ 0.028 0.476 $\pm$ 0.017 0.250 $\pm$ 0.009 0.100 $\pm$ 0.025 0.309 $\pm$ 0.019
0.05 -0.677 $\pm$ 0.039 0.084 $\pm$ 0.066 0.387 $\pm$ 0.002 0.359 $\pm$ 0.018 0.043 $\pm$ 0.031 0.372 $\pm$ 0.018 -0.215 $\pm$ 0.008 0.001 $\pm$ 0.030 0.044 $\pm$ 0.027
nRMSE $(\downarrow)$
0.01 16.858 $\pm$ 0.048 14.712 $\pm$ 0.058 12.747 $\pm$ 0.069 18.752 $\pm$ 0.092 15.937 $\pm$ 0.133 12.861 $\pm$ 0.094 15.070 $\pm$ 0.107 20.346 $\pm$ 0.203 15.910 $\pm$ 0.101
0.03 22.348 $\pm$ 0.391 15.658 $\pm$ 0.318 14.222 $\pm$ 0.053 19.085 $\pm$ 0.182 18.367 $\pm$ 0.349 14.551 $\pm$ 0.230 17.494 $\pm$ 0.107 21.187 $\pm$ 0.292 17.864 $\pm$ 0.240
0.05 31.452 $\pm$ 0.367 18.672 $\pm$ 0.675 15.524 $\pm$ 0.029 19.645 $\pm$ 0.280 20.783 $\pm$ 0.352 15.928 $\pm$ 0.236 22.266 $\pm$ 0.076 22.317 $\pm$ 0.338 20.823 $\pm$ 0.294
Table 29: MAE_FR_LP: Noise robustness analysis. Model performance under different noise intensities (0.01, 0.03, 0.05). Trait-wise $R^{2}$ (higher is better) and nRMSE (lower is better) are reported as mean $\pm$ standard deviation.
Noise cab cw cm LAI cp cbc car anth Average
$R^{2}$ $(\uparrow)$
0.01 0.583 $\pm$ 0.012 0.645 $\pm$ 0.024 0.787 $\pm$ 0.006 0.648 $\pm$ 0.014 0.667 $\pm$ 0.026 0.781 $\pm$ 0.007 0.584 $\pm$ 0.037 0.447 $\pm$ 0.031 0.643 $\pm$ 0.020
0.03 0.440 $\pm$ 0.028 0.536 $\pm$ 0.007 0.659 $\pm$ 0.009 0.530 $\pm$ 0.021 0.496 $\pm$ 0.035 0.647 $\pm$ 0.012 0.461 $\pm$ 0.029 0.264 $\pm$ 0.029 0.504 $\pm$ 0.021
0.05 0.242 $\pm$ 0.045 0.388 $\pm$ 0.023 0.477 $\pm$ 0.008 0.411 $\pm$ 0.014 0.276 $\pm$ 0.060 0.460 $\pm$ 0.009 0.258 $\pm$ 0.027 0.136 $\pm$ 0.073 0.331 $\pm$ 0.032
nRMSE $(\downarrow)$
0.01 15.677 $\pm$ 0.227 11.628 $\pm$ 0.388 9.141 $\pm$ 0.121 14.564 $\pm$ 0.290 12.265 $\pm$ 0.478 9.401 $\pm$ 0.147 13.025 $\pm$ 0.575 16.594 $\pm$ 0.462 12.787 $\pm$ 0.336
0.03 18.163 $\pm$ 0.461 13.302 $\pm$ 0.099 11.571 $\pm$ 0.146 16.823 $\pm$ 0.378 15.083 $\pm$ 0.526 11.947 $\pm$ 0.200 14.830 $\pm$ 0.406 19.154 $\pm$ 0.380 15.109 $\pm$ 0.325
0.05 21.131 $\pm$ 0.621 15.270 $\pm$ 0.291 14.346 $\pm$ 0.105 18.836 $\pm$ 0.219 18.078 $\pm$ 0.748 14.772 $\pm$ 0.123 17.400 $\pm$ 0.320 20.738 $\pm$ 0.879 17.571 $\pm$ 0.413
Table 30: MAE_FR_FT: Noise robustness analysis. Model performance under different noise intensities (0.01, 0.03, 0.05). Trait-wise $R^{2}$ (higher is better) and nRMSE (lower is better) are reported as mean $\pm$ standard deviation.
Figure 15: Feature importance of MAE-based downstream regression. Results are shown for (top) linear probing (MAE_FR_LPMAE\_FR\_LP), (middle) fine-tuning the last block (MAE_FR_9B_FTMAE\_FR\_9B\_FT), and (bottom) full fine-tuning (MAE_FR_FTMAE\_FR\_FT). The blue lines indicate the importance scores across spectral bands, while the orange line shows a reference of a vegetation spectra.